OpenStudy (anonymous):

HELP PLEASE? WILL MEDAL If g(x) = 2 (x-4), find the value of x if g(x) = 20 A) 32 B) 12 C) 14 D) 10 So how do I figure out what I plug in to get the answer?

Nnesha (nnesha):

replace g(x) with 20

OpenStudy (anonymous):

Alright, that's what I thought, I just wanted to make sure. I'm not really all that confident in myself when it comes to math lol

Nnesha (nnesha):

\[\huge\rm \color{Red}{g(x)}=2(x-4)\]\[\huge\rm \color{Red}{20}=2(x-4)\] now solve for x

OpenStudy (anonymous):

x = 14

Nnesha (nnesha):

thats mean you're pretty good at math :=)

Nnesha (nnesha):

yes right you can check ur answer substitute x for 14 \[20=2(\color{red}{14}-4)\] both sides are equal ? go.od to go!

OpenStudy (anonymous):

Thank you so much, I really appreciate it

Nnesha (nnesha):

np :=) good work! keep it up!